# Flamelet Connection to Turbulence Kinetic Energy Dissipation Rate

### v1

@@ Abstract
An analysis takes the variable value for turbulence kinetic energy dissipation rate $\epsilon$ as it might appear from a turbulent combustion computation using either Reynolds-averaged Navier-Stokes (RANS) or large-eddy simulation (LES) and relates it to both viscous dissipation rate and turbulence kinetic energy at the Kolmogorov scale. The imposed strain rate and vorticity on these smallest eddies are readily and uniquely determined from knowledge of that kinetic energy and viscous dissipation rate. Thus, a given value of $\epsilon$ at a specific time and location determines the two mechanical constraints (vorticity and strain rate) on the inflow to the flamelet. It is also shown how $\epsilon$ affects the sign of the Laplacian of pressure, which must be negative to allow the existence of the flamelet. Using several different flamelet models, with and without vorticity and with and without differential mass transport, different results for maximum flamelet temperature, integrated flamelet burning rate, and stoichiometric flamelet scalar dissipation rate are obtained. For a given $\epsilon$ value, flamelet models that do not consider vorticity and differential diffusion produce substantial errors in the information to be provided to the resolved or filtered scales in a turbulent combustion computation.
1 Introduction
@@
The major method for energy conversion for high levels of mechanical power and heating continues to be combustion with high mass fluxes. Inherently, the high mass-flow rate involves turbulent flow. Consequently, many physical length and time scales emerge and present challenges for both computational and experimental analyses. Large-eddy simulations (LES) are employed for computations where the smallest scales cannot be resolved. The smaller scales are filtered via integration over a window size commensurate with the computational mesh size, thereby allowing affordable computations. Consequently, the essential, rate-controlling, physical and chemical processes that occur on scales smaller than the filter size must be described by flamelet closure models. The flamelet models must handle multi-species, multi-step oxidation kinetics without requiring small time steps during the solution of the resolved-scale fluid dynamics. Thus, savings of computational resources can be huge compared to direct numerical simulation. The focus here is on the types of turbulent flames found in the shear-driven flows of practical combustors. Thereby, we address flamelets as originally described by Williams (1975), namely “highly sheared small diffusion flamelets” and “forming a turbulent flame brush which appears on the average to fill” the flow domain.
The goal here is to establish useful connections of Reynolds-averaged Navier-Stokes (RANS) and large-eddy simulation (LES) frameworks. The formulation developed here becomes a candidate for relating the resolved scale of RANS or LES to the subgrid scale for the flamelet. For the LES, we especially aim for detached-eddy simulations where, like RANS, resolutions of turbulence kinetic energy $k$ and turbulence kinetic energy dissipation rate $\epsilon$ are determined to describe the flow below the filtered length scale.
@@
The brief literature review of the flamelet models is presented in Section 2. A brief overview of key relations for the rotational flamelet model is given in Section 3. The description of the velocity divergence and curl in the rotating flamelet reference frame is discussed in Section 4 while the viscous dissipation, enstrophy, and kinetic energy in that frame are discussed in Sections 5 and 6. Scaling of velocity gradients between the resolved scale and flamelet level are analyzed in Section 7; specific use of and connection of the flamelet inflow with turbulence kinetic energy dissipation rate is addressed in Section 8. Results of calculations for several flamelet model forms are discussed in Section XX and Conclusions are given in Section XXX.
2 Flamelet Closure Models
@@
It is critical to understand the laminar mixing and combustion that commonly occur within the smallest turbulent eddies. These laminar flamelet sub-domains experience significant strain of all types: shear, tensile, and compressive. Notable works exist here for either counterflows with only normal strain or simple vortex structures in two-dimensions or axisymmetry and often with a constant-density approximation. See Liñán (1974), Marble (1985), Karagozian & Marble (1986)), Cetegen & Sirignano (1988), Cetegen & Sirignano (1990), Peters (2000), and Pierce & Moin (2004). An interesting review of the early flamelet theory is given by Williams (2000). Flamelet studies have focused on either premixed or nonpremixed flames; a unifying approach to premixed, nonpremixed, and multi-branched flames is needed.
@@
Vorticity interaction with the flame has not been directly considered in most models (Liñán, 1974; Peters, 2000; Williams, 2000; Pierce & Moin, 2004). Two-dimensional planar and axisymmetric counterflow configurations are generally the foundation for a flamelet model. Local conversion to a coordinate system based on the principal strain-rate directions can provide the counterflow configuration in a general flow. Furthermore, the quasi-steady counterflow can be analyzed by ordinary differential equations because the dependence on the transverse coordinate is either constant or linear, depending on the variable. The mixture fraction has been used widely as an independent variable to display non-premixed flamelet scalar variations; this cannot be useful for premixed flames. Sirignano (2021a) has shown that any conserved scalar can serve well as an independent variable to present scalar results for nonpremixed and multi-branched flamelets.
Hamins et al. (1985) have shown through experiments and asymptotic analysis that a partially premixed fuel-lean flame and a diffusion flame can co-exist in a counterflow with opposing streams of heptane vapor and methane-oxygen-nitrogen mixture. Thus, a need exists for flamelet theory to address both premixed and non-premixed flames. Recently, Rajamanickam et al. (2019) provided an interesting three-dimensional triple-flame analysis, describing the effect of imposed normal strain on a multibranched flame. They did not consider shear strain, but it was followed by the work of Sirignano (2021b) where both shear strain and normal strain were considered. The classical counterflow treatment (Bilger, 1976; Liñán, 1974; Peters, 2000) has two opposing streams, fuel or fuel plus a chemically inert gas and oxidizer or oxidizer plus an inert gas. They considered uniform density. That critical assumption was relaxed by (Sirignano, 2019) for reacting flows and heated flows. Sirignano (2021a, b) with one-step kinetics and López-Cámara et al. (2020) with detailed kinetics address situations where the inflowing streams consist of a combustible mixture of fuel and oxidizer, thereby allowing another flame or two besides the simple diffusion flame to co-exist. Sirignano (2021a) provides a counterflow analysis that shows the various permissible flame configurations pertaining to the specific compositions of the inflowing streams: (i) three flames including fuel-lean partially premixed, nonpremixed (i.e., diffusion-controlled), and fuel-rich partially premixed; (ii) nonpremixed and fuel-rich partially premixed; (iii) fuel-lean partially premixed and nonpremixed; (iv) nonpremixed; and (v) premixed. López-Cámara et al. (2020) extended the counterflow analysis to consider detailed kinetics for methane-oxygen detailed chemical kinetics and confirmed that combinations of premixed and non-premixed flames could exist in a multi-flame counterflow.
The effects of both normal strain rate and shear strain rate should be determined. Mixing and combustion in three-dimensional flows need to be studied with imposed normal strain and shear strain and therein imposed vorticity. Several new flamelet models have been developed for use in sub-grid closure modeling. They would be coupled with the resolved flow description for turbulent combustion. These models differ from current models in several critical ways. (i) Non-premixed flames, premixed flames, or multi-branched flame structures are determined rather than prescribed. (ii) The effects of shear strain and vorticity on the flames are determined. (iii) The strain rates and vorticity applied at the sub-grid level are directly determined from the resolved-scale strain rates and vorticity without the use of a contrived progress variable. (iv) The flamelet model is three-dimensional without need for assuming axisymmetry or planar geometry. (v) The effect of variable density is addressed in the flamelet model. Solutions to the Navier-Stokes equations and the associated scalar equations governing the flamelet model are obtained without the boundary-layer approximation.
@@
By appropriate coordinate transformation, a similar solution is found for each flamelet model, reducing it to system of ordinary differential equations (ODEs). Sirignano (2022c) considers a rotational flamelet with uniform vorticity over the small flamelet domain; t2022che three-dimensional counterflow has one compressive-normal-strain direction aligned with scalar gradients and two extensional-normal-strain directions. Sirignano (2022a) considers a stretched vortex tube with inward swirl created by two compressive-strain directions and the extensional direction aligned with the vorticity vector. Here, we have an analog to the classical Burgers stretched vortex (Burgers, 1948; Rott, 1958). In both models, vorticity is shown to create a centrifugal force on the sub-grid counterflow that modifies the molecular transport rates and burning rate. This has consequence on the flow rate through the flame zone and the burning rate. It thereby modifies flammability limits. Another model has been introduced Sirignano (2022b) where shear strain is explicitly represented in the $x,y$ plane creating a vortex layer with counterflow in the $y,z$ plane producing vortex stretching. This model is an analog to the Burgers vortex-layer model (Burgers, 1951; Batchelor, 1967; Robinson & Saffman, 1984). In the rotational flamelet and the stretched-vortex-tube three-dimensional models, the imposed vorticity had consequence on the scalar fields and the burning rate through the centrifugal acceleration. However, here, the vortex-layer model is asymptotically two-dimensional with no centrifugal effect, and a resulting one-way coupling between the scalar field and the vorticity; vorticity does not affect scalar profiles while the scalars can have some affect on vorticity.
The analytical frameworks allow for multi-step, detailed kinetics although the presented calculations are limited to one-step kinetics. Hellwig et al. (2023a) has recently extended the analysis to include detailed kinetics and transport and variable thermophysical properties with a more complete examination of flammability limits. The quasi-steady assumption commonly used in previous flamelet models is maintained here for the flamelet. Premixed flames can only exist with orders-of-magnitude larger values of $Da$ compared to diffusion flamelets or multibranched flames. So, premixed flames will be less likely to survive extinction in a turbulent situation with high strain rates. The partially premixed flames in a multi-branched flame structure are driven by heat flux from the diffusion flame; thereby, they survive at lower values of Damköhler number $Da$. Hellwig et al. (2023b) examined the scalar dissipation-rate prediction of the rotational-flamelet model in the case where Fickian diffusion at the same rate is assumed for all species, a condition used in the Progress Variable Method, which neglected vorticity and implicitly used only a compressive normal strain rate. The rotational flamelet has the direct input of both strain rate and vorticity. For both models, it is found that a given maximum temperature or reaction rate correlates with a specific scalar dissipation rate (SDR) in the reaction zone. However, different imposed values of compressive strain rate and vorticity impose that same value of SDR.
@@
Scaling laws have been presented (Sirignano, 2022c) for relating strain rates and vorticity at the sub-grid level to those quantities at the resolved-flow level for coupling with large-eddy simulations or the time-averaged mean-flow level for Reynolds-averaged flows. The time-averaged behavior of a simple turbulent flow is resolved with coupling to the rotational flamelet model. Specifically, a two-dimensional, multicomponent, time-averaged planar shear layer with variable density and energy release is employed using a mixing-length description for the eddy viscosity.
@@
Useful background information relevant to turbulent nonpremixed combustion is given by Elghobashi & Pun (1974); Elghobashi (1977); Nomura & Elghobashi (1992, 1993); Boratav et al. (1996, 1998); Ashurst et al. (1987); Hamlington et al. (2008). Background for premixed turbulent combustion is found with Driscoll et al. (2020); Steinberg et al. (2021); Steinberg & Driscoll (2009); Hamlington et al. (2010); Swaminathan & Grout (2006).
@@
Here, we address the use of turbulence kinetic energy dissipation rate $\epsilon$ to connect the RANS resolved-scale or LES filtered-scale for turbulent non-premixed combustion to the sub-grid falmelet model. $\epsilon$ is a well establish parameter in turbulence theory (Jones & Launder, 1972; Pope, 2000)and will be used as tracking variable to determine the parameters (vorticity and strain rate) at the smallest scale of turbulence that are constraints for the flamelet model. Specifically, we work here with the rotational flamelet model presented by Sirignano (2022a), Sirignano (2022c), Hellwig et al. (2023a).

### v2

@@ Abstract
The turbulence kinetic energy dissipation rate $\epsilon$, from a turbulent combustion computation using either Reynolds-averaged Navier-Stokes (RANS) or large-eddy simulation (LES), is proposed for closure with a sub-grid non-premixed flamelet model. The intentions are to avoid the creation of artificial tracking or progress variables and to relate accurately the physics of turbulent non-premixed combustion at the resolved length scales to the small-scale physics where the mixing and chemical reactions occur. The analysis addresses the relations between $\epsilon$ and the strain rate, vorticity, viscous dissipation rate, scalar gradients, scalar dissipation rate, and burning rate at the smallest turbulence length scales where diffusion-controlled burning is faster than at larger length scales and thereby dominant. The imposed strain rate and vorticity on these smallest eddies are determined from the kinetic energy dissipation rate. Thus, an $\epsilon$ value at a specific time and location determines the two mechanical constraints (vorticity and strain rate) on the inflow to the counterflow flamelet. $\epsilon$ affects the sign of the Laplacian of pressure, which must be negative to allow the existence of the counterflow. Using different flamelet models, with and without vorticity, different results for maximum flamelet temperature, integrated flamelet burning rate, and maximum flamelet scalar dissipation rate are obtained. Flamelet models that consider the centrifugal effect of vorticity produce substantial enhancements in the accuracy and completeness of information for a turbulent combustion computation. $\epsilon$ may be used as a tracking variable that connects the sub-grid flamelet model to resolved-scale RANS or LES computations. ARTICLE WORD COUNT: 8060 words.
keywords:
Flamelet model; closure tracking variable;
@@
Large-eddy simulations (LES) or Reynolds-averaged Navier-Stokes (RANS) are used for turbulent combustion computations where the smallest scales cannot be resolved. The smaller scales are filtered via integration over a window size commensurate with the computational mesh size, thereby allowing affordable computations. The essential, rate-controlling, physical and chemical processes that occur on scales smaller than the filter size must be described by flamelet closure models. The flamelet models must handle multi-species, multi-step oxidation kinetics without requiring small time steps during the solution of the resolved-scale fluid dynamics. The focus here is on the types of turbulent flames found in the shear-driven flows of practical combustors. Thereby, we address flamelets as originally described by Williams [Williams1975], namely “highly sheared small diffusion flamelets” and “forming a turbulent flame brush which appears on the average to fill” the flow domain.
The use of the turbulence kinetic energy dissipation rates to connect large-length-scale behavior to the smallest scale behavior is well established for the mechanics of turbulent flow [Pope2000]. It has been used for many decades to relate resolved-scale behavior in Reynolds-averaged Navier-Stokes (RANS) solvers and large-eddy simulations (LES) to sub-grid behavior. Here, with a clear connection between scalar gradients on the smallest scale and the resolved-scale mechanics, we propose a new method to relate the small scales where diffusion-controlled combustion occurs to the resolved-scale flow. Focus will be placed on the two-way coupling, including the feedback of the heat release rate from the small-length scales to the resolved scale.
In its most basic sense, the flamelet model quantifies the unresolved fluid-chemistry interaction occurring at each RANS or LES node or cell. The fluid-chemistry interaction will be determined by a set of governing equations which are subject to two types of constraints: 1) mechanical constraints, indicative of velocity gradients and thus residence time; and 2) thermodynamic constraints, which are indicative of mixture, temperature, and pressure. These constraints on the flamelet model must be a function of known RANS or LES quantities for the overall CFD framework to close the system. In this way, the coupling between RANS or LES and the flamelet model parallels the classical closure approach used to relate kinetic energy on the averaged or resolved scale to the sub-grid dissipation of that energy. We specifically target the mechanical constraints of the Rotational Flamelet Model (RFM) [Sirignano2022a, Sirignano2022b, Hellwig2025a, Hellwig2025b]; however, the procedure laid out in this work can also be applied to other flamelet models. This work does not address a new coupling for thermodynamic or chemical constraints.
The formulation here does not input scalar dissipation rate or progress variable into the subgrid flamelet model and does not use a probability density function but rather provides an alternative coupling procedure with a differing foundation. For LES, we especially aim for detached-eddy simulations where, like RANS, resolutions of turbulence kinetic energy $k$ and turbulence kinetic energy dissipation rate $\epsilon$ are determined to describe the flow below the filtered length scale. The coupling for the mechanical constraints established here is conditionally formulated on the ability of the RANS or LES code to quantify $\epsilon$.
The method builds upon a few established concepts in the science of turbulence. The cascade of turbulence kinetic energy through the length scales has mechanical energy being transferred from larger eddies to sequentially smaller eddies with insignificant dissipation until the range of smallest eddy sizes is reached. Although there is little turbulence kinetic energy in these smallest (Kolmogorov and near-Kolmogorov) scales, the viscous dissipation rate dominates there because they contain the largest velocity derivatives. Thereby, these scales also contain the largest values of vorticity (or enstrophy) and normal strain rates in the spectrum. Since viscosity is the diffusion of momentum or vorticity, we expect, for turbulent flows with near-unity values of Prandtl number and Schmidt number, scalar gradients and mixing rates for heat and mass will be dominate on those same length scales, where kinetic energy dissipation occurs. Furthermore, since the smallest eddies have low values of kinetic energy, compared to larger length scales in the spectrum (as shown by Figure 6.22 of Pope 2000 [Pope2000]), their impact on eddy diffusivity or flow curvature on larger scales is not significant. However, molecular mixing and diffusion-controlled combustion will dominate on this scale. On average, we assume the distance travelled by the gaseous material in the turbulence cascade from the end of the computational resolution to the dissipation completion is small. However, since the distance between two material points can decrease due to sever straining, scalar gradients will increase as length scale decreases.
Some valuable guidance on the relation between scalar mixing and viscous dissipation is provided by the work of Yeung, Donvis, and Sreenivasan (YDS) [yeung_high-reynolds-number_2005]. Following them, $\delta u$ is the velocity increment over a distance $\delta r$ in the same direction as the velocity vector. $\delta\phi$ is the conserved scalar increment in that same direction and distance. $\phi$ is non-dimensional and $O(1)$ over the flow. From Equations (1, 3) and Figures (1, 3) of the YDS paper, $\left(\delta u\right)^{3}/\left(\epsilon\delta r\right)=4/5$, where $\epsilon$ is the viscosity-based turbulence kinetic energy dissipation rate. The scalar dissipation rate is $\chi=\left(3/2\right)\left(\delta\phi\right)^{2}\delta u/\delta r$. Based on this definition, we also expect $\chi=2D\left(\delta\phi/\delta r\right)^{2}$ where $D$ is the thermal or mass diffusivity. Now, setting $\delta u\delta r/\nu=1$ (for Kolmogorov scales where $\nu$ is the kinematic viscosity), we have $\epsilon/\nu=\left(5/4\right)\left(\delta u/\delta r\right)^{2}$ and $\delta u/\delta r=\left(2/3\right)\chi/\left(\delta\phi\right)^{2}$. Thus, $\chi=3(\delta\phi)^{2}\left(\epsilon/(5\nu)\right)^{\frac{1}{2}}$ where $\delta\phi$ is a nondimensional variable smaller than $O(1)$. Also, $\chi\sim\delta u/\delta r$. That is, strain rate determines scalar dissipation rate on this scale.
@@
A brief literature review of the Classical Flamelet Model (CFM) and its mechanical coupling procedures is presented in Section 1. An overview of the Rotational Flamelet Model (RFM) is presented in Section 2. An overview of key relations for the RFM is given in Section 3. The description of the velocity divergence and curl in the RFM rotating reference frame is discussed in Section 4 while the viscous dissipation, enstrophy, and kinetic energy in that frame are discussed in Sections 4 and 5. Scaling of velocity gradients between the resolved scale and flamelet scale are analyzed in Section 6; specific use of and connection of the flamelet inflow with turbulence kinetic energy dissipation rate is addressed alongside results of calculations for several flamelet models in Section 7 and Conclusions are given in Section 8. While the science and key relations are provided for two-way flamelet-flow connection, the computational application with a computation for a turbulent combustion field is left for future study.
1 Classical Flamelet Model and Coupling Procedure for Mechanical Constraints
@@
It is critical to understand the laminar mixing and combustion that commonly occur within the smallest turbulent eddies. These laminar flamelet sub-domains experience significant strain of all types: shear, tensile, and compressive. Notable works exist here for either counterflows with only normal strain or simple vortex structures without stretching in two-dimensions or axisymmetry and sometimes with a constant-density approximation implied through an assumption of uniform nomral strain rate. See [Peters, Pierce, Linan, Marble, Karagozian, Cetegen1, Cetegen2]. An interesting review of the early flamelet theory is given by [Williams2000] and a recent review is given by [DOMINGO2025]. Flamelet studies have generally focused on either premixed or non-premixed flames, although some recent works have adopted a unifying approach to premixed, non-premixed, and multi-branched flames, referred to as multi-modal combustion.
@@
Vorticity interaction with the flame has not been directly considered in most models [Peters, Pierce, Linan, Williams2000]. Two-dimensional planar and axisymmetric counterflow configurations are generally the foundation for a flamelet model. Local conversion to a coordinate system based on the principal strain-rate directions typically provides the counterflow configuration. Furthermore, the quasi-steady counterflow can be analyzed by ordinary differential equations because the dependence on the transverse coordinate is either constant or linear, depending on the variable. The mixture fraction has been used widely as an independent variable to display non-premixed flamelet scalar variations; this cannot be useful for premixed flames. It has been shown by [Sirignano2021a] that any conserved scalar can serve well as an independent variable to present scalar results for nonpremixed and multi-branched flamelets.
The classical counterflow treatment [Linan, Peters, Bilger], hereafter refereed to as the “Classical Flamelet Model” (CFM), solves the flamelet equations in mixture fraction space, yielding laminar reactive scalars $\psi_{i}=\psi_{i}(Z,\chi)$. Once these are obtained, a coupling procedure is executed whereby resolved scale RANS or LES quantities indicate which of the many $\chi$ (mechanical constraint) profiles to use for the given resolved-scale node or cell. A similar coupling must be given for the thermodynamic state, but we do not discuss this in detail here. The mechanical coupling is typically performed in one of two ways using the presumed PDF approach: 1) use $\chi$ directly in a PDF (This method involves only the two stable branches of the S-curve and does not involve the unstable branch as possible flamelet states); or 2) bijectivley map $\chi$ to a monotonic Flamelet Progress Variable (FPV) $C$ [Pierce] and use that variable along with a presumed PDF approach (this method includes all branches of the S-curve as realizable flamelet states).
In a presumed PDF approach, the Favre mean $\widetilde{\chi}$ is an input to the PDF and must be modeled on the resolved scale. This is commonly done as $\widetilde{\chi}=C_{\chi}(\widetilde{\epsilon}/\widetilde{k})\widetilde{Z^{\prime\prime 2}}$ where $C_{\chi}$ is the proportionality constant between turbulence and scalar time-scales. $C_{\chi}$ is typically set to 2.0; however, this value was determined specifically with an inert jet of methane [janicka_prediction_1982]. A modeled equation can also be used, see [Elghobashi1977, ElghobashiLaunder1983, NewmanLaunderLumley1981, BorghiDutoya1979].
The purpose of the PDF for $\chi$ is to couple the LES or RANS quantities to the flamelet quantities. Flamelets exist on length scales smaller than the LES filter scale and below the mean produced by RANS; thus, the flamelet should not be directly subjected to the LES or RANS quantities, but rather to the instantaneous flow field that is filtered or averaged away by LES and RANS. Of course, the actual instantaneous flow field is unknown. So, a model of the turbulent fluctuations is used. The log-normal PDF $P(\chi)$ is one such model which, given a mean $\widetilde{\chi}$, describes the probability of instantaneous scalar dissipation rate occurring in the range $0<\chi<\infty$. Since the flamelet equations are solved for the same range of $\chi$, convolution between $\psi_{i}(Z,\chi)$ and $P(\chi)$ is claimed to characterize the fluctuations and intermittency of instantaneous small-scale eddies and their effect on combustion. The log-normal nature of $P(\chi)$ stems from Kolmogorov’s third hypothesis and has been shown to be in good agreement with experiments and numerical data [SreenivasanAntoniaDanh1977, SreenivasanAntonia1977, Kerstein1984, Dahm1989] provided the length scale observed was between $\kappa<<r<<L$ where $\kappa$ is the Kolmogorov length scale and L is the integral length scale. Sreenivasan et al. [SreenivasanAntoniaDanh1977] observe lognormality of the PDF between $5\kappa<r<150\kappa$. Other authors have used a delta function for $P(\chi)$ centered on $\widetilde{\chi}$ which completely ignores fluctuations. Sanders and Lamers [sanders_modeling_1994] match experiments better in the prediction of turbulent diffusion flame liftoff height, using a Gaussian PDF for strain rate of the smallest length scales, rather than a log normal PDF for scalar dissipation rate. The mean and variance of that strain rate PDF are proportional to the square root of the turbulence kinetic energy dissipation rate in their RANS formulation. That dependence relates to our analysis and results in the following sections.
Regardless of the PDF used, the present authors are concerned about a mechanical coupling using $\chi$ for two related reasons. First, it requires a presumed profile $\chi(Z)$ to solve the CFM equations which invokes multiple simplifying assumptions and thus far has not included the effect of Kolmogorov vorticity at the smallest scales [Sirignano2022a, Sirignano2022b, Hellwig2025a, Hellwig2025b]. Furthermore, the definition of scalar dissipation rate of a conserved scalar requires equality among the various species diffusivities. We do not wish to invoke these various assumptions. Second, we hypothesize that the diffusion layer held by a Kolmogorov eddy and other small scale eddies see the influence of mechanical constraints as far-field conditions. The flame-local structure is then determined through governing equations subjected to three far-field strain rates and vorticity. Hellwig et al. [Hellwig2025a] shows that inclusion of these four parameters creates a multi-valued correspondence between $S^{*}$ and $\chi_{st}$ whereas the assumed profile $\chi(Z)$ produces a one-to-one correspondence. Therefore, it is preferred to use a formulation considering the asymptotic strain rate, $S^{*}$, as well as the other two transverse strain rates and vorticity, rather than an assumed profile for $\chi$ based on $S^{*}$ alone. We seek a coupling procedure that can quantify these mechanical constraints. Furthermore, the use of a PDF for $\chi$ introduces an inconsistency with the scalar dissipation rate term in the resolved-scale equation for scalar variation (the variance of the mixture fraction in the CFM approach), namely, $\widetilde{\chi}=C_{\chi}(\widetilde{\epsilon}/\widetilde{k})\widetilde{Z^{\prime\prime 2}}$ where the characteristic time for that rate is assumed to be proportional to the characteristic time for energy dissipation rate. There are implicitly two scalar dissipation rate definitions involved.
These concerns are exacerbated when adopting a progress variable (FPV) framework. In addition to requiring the same assumptions about scalar dissipation rate profiles and the modeling of the scalar dissipation rate on the resolved scale, the FPV approach introduces a more fundamental issue: a loss of the mechanical constraint fidelity. Walsh et al. [walsh_turbulent_2025] found that once flamelet solutions are mapped from scalar dissipation rate (i.e., strain rate) to progress variable the, intrinsic relationship between strain rate and flamelet structure is no longer preserved in the resolved-scale solution. This decoupling arises because the progress variable evolves according to its own transport equation at the resolved scale, independent of the local strain rate, despite the fact that the strain rate governs the underlying flamelet structure in the pretabulated solutions. As a consequence, the resolved-scale model cannot retain the correlation between strain rate and the progress variable values that exists at the subgrid flamelet level. It has been shown [walsh_performance_2025] that this decoupling leads to the preferred selection of equilibrium flamelet (low strain solutions) solutions in regions of high strain, resulting in nonphysical heat release rates. So here, the mechanical constraint becomes non-strain rate informed.
The theory presented in this paper is based on the hypothesis that the smallest eddies, those on or around the Kolmogorov scale and which have the highest strain rates and shortest time scales, have leading order effects on combustion since they produce the highest heat release rates. Furthermore, these eddies on the near-Kolmogorov scales are the only eddies which are laminar and thus the only sensible location for a laminar flamelet. Therefore, the goal of this work is to derive a coupling procedure which approximates the strain rate and enstrophy $\omega^{2}$ magnitude of these particular eddies. We do not attempt to predict stochastically the variation in strain rate due to intermittency. Instead, we take a spatial average over a Kolmogorov volume $\kappa^{3}$ to approximate the average mechanical constraint experienced by that Kolmogorov eddy. A temporal average needed to account for intermittency will be applied through a coefficient. In other words, the fluctuations of Kolmogorov eddies will be averaged out.
We present here an alternative coupling procedure which connects sub-grid strain rate and thereby sub-grid scalar behavior directly to the turbulence kinetic energy dissipation rate $\epsilon$. Specifically, we connect the sub-grid scalar dissipation rate to the resolved scale whereby $\epsilon$ is used to quantify the strain rate $S^{*}$ and the vorticity magnitude $\omega$, that a laminar counterflow flamelet existing on the Kolmogorov scale, would experience without the use of a PDF approach. These asymptotic quantities are grounded firmly in turbulence cascade theory and may thus offer a more physically-sound coupling of the mechanical constraints between scales for the specific viewpoint of the flamelet model as a Kolmogorov-scale model. Although the approach is developed with the RFM in mind, as shown in Table 1, it may also be applied to the CFM approach, as will become evident in Section 7.
2 Rotational Flamelet Model
@@
The Rotational Flamelet Model (RFM) is a subgrid, Kolmogorov-scale flamelet model originally created by Sirignano [Sirignano2022a]. It uses a 3D, non-Newtonian reference frame rotating at half of the vorticity magnitude such that shear strain is removed from the inflows; i.e., three principal strain rates exist. The major tensile strain rate is $S_{1}^{*}=S_{1}S^{*}$, the major compressive strain rate is $S^{*}$, and the intermediate strain rate, which may be tensile or compressive, is $S_{2}^{*}=S_{2}S^{*}$. Scalar gradients are aligned with the major compressive strain rate direction. Vorticity and the stretching of vorticity align with the major extensional direction. Through the coordinate transformation, a similar solution is found for the governing equations, reducing it to a system of ordinary differential equations (ODEs). The vorticity is uniform across the small flamelet domain. This is justified because vorticity maximizes towards the center of the eddy creating a zero gradient of vorticity locally. Sirignano considered both a situation where the intermediate strain rate was tensile [Sirignano2022a] where a stretched vortex sheet or layer is created, and a stretched vortex tube situation with inward swirl created by two compressive-strain directions and the extensional direction aligned with the vorticity vector [Sirignano2022b]. This tube presented an analog to the classical Burgers stretched vortex [Burgers1948, Rott]. In both models, vorticity is shown to create a centrifugal force on the sub-grid counterflow that modifies the molecular transport rates and burning rate. This has consequence on the flow rate through the flame zone and the burning rate. It thereby modifies flammability limits. These initial demonstrations of the RFM were done with Fickian diffusion, unity Lewis numbers, and one-step chemical kinetics.
Hellwig et al. [Hellwig2025b] improved the rotational flamelet model by including detailed kinetics, multicomponent transport, and variable thermophysical properties with a more complete examination of flammability limits. The quasi-steady assumption commonly used in previous flamelet models was retained here for the flamelet. Hellwig et al. [Hellwig2025a] later substituted a Fickian diffusion formulation, setting the diffusivities of all species equal to the thermal diffusivity in order to define scalar dissipation rate in the RFM and showed that vorticity and varying transverse strain rates create a multi-valued correspondence between the asymptotic strain rate $S^{*}$ and the flame-local scalar dissipation rate. This work was conducted because the scalar dissipation rate is given importance in the CFM and FPV methods, which neglected vorticity 3D effects. The RFM has the direct input of both strain rate and vorticity. For both models, a given maximum temperature or burning rate correlates with a specific scalar dissipation rate (SDR) in the reaction zone. However, different imposed values of compressive strain rate and vorticity impose that same value of SDR. Both [Hellwig2025a, Hellwig2025b] showed fundamental results consistent with the original works of Sirignano [Sirignano2022a, Sirignano2022b], in that the combination of varying transverse strain rates and vorticity modify the local flame structure and burning rate.
@@
Scaling laws have been presented by [Sirignano2022a] for relating strain rates and vorticity at the sub-grid level to those quantities at the resolved-flow level for coupling with large-eddy simulations or the time-averaged mean-flow level for Reynolds-averaged flows. The time-averaged behavior of a simple turbulent flow is resolved with coupling to the RFM. Specifically, a two-dimensional, multicomponent, time-averaged planar shear layer with variable density and energy release is employed using a mixing-length description for the eddy viscosity.
@@
Useful background information relevant to turbulent non-premixed combustion is given by References [Elghobashi1977, Nomura1992, Elghobashi1974, Nomura1993, Boratav1996, Boratav1998, Ashurstetal, Dahm]. Background for premixed turbulent combustion is found with References [Driscoll, Steinbergetal, Steinberg, Hamlington, Swaminathan].
@@
Here, we address the use of turbulence kinetic energy dissipation rate $\epsilon$ to connect the RANS resolved-scale or LES filtered-scale for turbulent, globally non-premixed combustion to the sub-grid flamelet model. The quantity $\epsilon$ is a well-established parameter in turbulence theory [JonesLaunder, Pope2000] and will be used as a coupling variable to determine the parameters (vorticity and strain rate) at the smallest scale of turbulence that are constraints for the flamelet model. Specifically, we address here the inputs to the RFM [Sirignano2022a, Hellwig2025a, Hellwig2025b, Sirignano2022b]. We also make comparisons with CFM results generated by Flamemaster [flamemaster], which are based on references [Peters, Pitsch1, Pitsch2].